OpenStudy (anonymous):

To simplify \[5\div2+\sqrt{6}\] I think you would multiply the top and bottom by \[2-\sqrt{6}\] but I'm not positive. Can someone help?

OpenStudy (anonymous):

you could that

OpenStudy (unklerhaukus):

is that \[\frac5{2+\sqrt6}\]? if so , yes, multiply by \[\frac{2-\sqrt 6}{2-\sqrt 6}\]

OpenStudy (unklerhaukus):

This technique is called rationalising the denominator , it will get all those horrible surds out of the denominator

OpenStudy (anonymous):

would it them become just 6?

OpenStudy (anonymous):

then*

OpenStudy (unklerhaukus):

no